# footnotetext: $\dagger$ First authors

### v1

@@ B.1 Appendix: Fukumizu Approach
The four quadrants of 28 are equivalent to equations 32, 35, 38, and 43 respectively.
@@ B.3 $\bm{F}\quad$ Diagonalization
Under assumptions of full-rank 4, the eigendecomposition of $\mathbf{F}=\mathbf{P}\bm{\Lambda}\mathbf{P}^{T}$ where
@@ B.4 Solution Unequal-Input-Output
Under the assumptions of whitened inputs, 4, lambda-balanced weights 4, no bottleneck 4, and full rank 4, the temporal dynamics of ${\bf Q}{\bf Q}^{T}$ are
@@ C.2 Dynamics of the representation from the Lazy to the Rich Regime
The lazy and rich regimes are defined by the dynamics of the NTK of the network. Lazy learning occurs when the NTK is constant, rich learning occurs when it is not. (Farrell et al. (2023b))
@@ C.2.1 Lambda-balanced singular value
Under a $\lambda$-Balanced initialization 4, if the network function $\bm{W}_{2}\bm{W}_{1}(t)=\bm{U}(t)\bm{S}(t)\bm{V}^{T}(t)$ is full-rank 4 and we define $\bm{S_{\lambda}}(t)=\sqrt{\bm{S}^{2}(t)+\frac{\lambda^{2}}{4}\mathbf{I}}$. , then we can recover the parameters $\bm{W}_{2}(t)=\bm{U}(t)\bm{S}_{2}(t)\bm{R}^{T}(t)$, $\bm{W}_{1}(t)=\bm{R}(t)\bm{S}_{1}(t)\bm{V}^{T}(t)$ up to time-dependent orthogonal transformation $\bm{R}$(t) of size $N_{h}\times N_{h}$, where
@@
By properties of Singular Value Decomposition, we can write $\bm{W}_{2}=\bm{U}\bm{S}_{2}\bm{R}^{T},\bm{W}_{1}=\bm{R}\bm{S}_{1}\bm{V}^{T}$, where $\bm{R}$ is an orthonormal matrix and $\bm{S}_{2},\bm{S}_{1}$ are diagonal (possibly rectangular) matrices.
The Balanced property states that $\bm{W}_{2}^{T}\bm{W}_{2}-\bm{W}_{1}\bm{W}_{1}^{T}=\lambda\mathbf{I}$. We know this holds for any $t$ since this is a conserved quantity in linear networks.
@@ C.2.4 NTK movement
The theorem above demonstrates that as $|\lambda|\to\infty$, the NTK of a $\lambda$-Balanced network remains constant. This indicates that the network operates in the lazy regime throughout all training steps. This finding is significant as it highlights the impact of weight initialization on learning regimes.
@@ C.3 Representation robustness and sensitivity to noise
As derived in (Braun et al., 2024), the expected mean squared error under additive, independent and identically distributed input noise with mean $\mu=0$ and variance $\sigma_{\bf x}^{2}$ is
$\left\langle\frac{1}{2P}\sum_{i=1}^{P}||{\bf W}_{2}{\bf W}_{1}\left({\bf x}_{\bf x}+\xi_{i}\right)-{\bf y}_{i}||_{2}^{2}\right\rangle_{\xi_{\bf x}}=\sigma_{\bf x}^{2}||{\bf W}_{2}{\bf W}_{1}||_{F}^{2}+c,$ (170)
@@
As derived in (Braun et al., 2024), under the assumption of whitened inputs (Assumption 4), in the case of additive parameter noise with $\mu=0$ and variance $\sigma_{\bf W}^{2}$, the expected mean squared error is
@@ C.4 Effect of the architecture from the lazy to the Rich Regime
As $\lambda$ increases, different parts of the network exhibit distinct learning behaviors: some components adapt quickly and converge exponentially with lambda, while others are constrained by the singular values of the network, resulting in slower adaptation.
@@
The first term exhibits exponential decay with rate $\lambda$, approaching zero as time progresses:
@@
The second term also decays, but at a rate governed by the singular values $\tilde{\bm{S}}$, as $\lambda$ tends to infinity:
@@
Since
$\lambda_{\perp}-\sqrt{\tilde{\bm{S}}^{2}+\frac{\lambda^{2}}{4}}\mathbf{I}>0,$
we have
$\lim_{\lambda\to\infty}\left(\lambda_{\perp}-\sqrt{\tilde{\bm{S}}^{2}+\frac{\lambda^{2}}{4}}\mathbf{I}\right)=\tilde{\bm{S}}.$
Thus, as $\lambda$ increases, the convergence rate slows for certain parts of the network (those governed by larger singular values), while other components continue to learn more quickly. This explains the delay observed in the delayed-rich regime. ∎

### v2

@@ B.1 Appendix: Fukumizu Approach
The four quadrants of 27 are equivalent to equations 31, 34, 37, and 42 respectively.
@@ B.3 $\bm{F}\quad$ Diagonalization
The eigendecomposition of $\mathbf{F}=\mathbf{P}\bm{\Lambda}\mathbf{P}^{T}$ where
@@ B.4 Solution Unequal-Input-Output
Under the assumptions of whitened inputs, 4, lambda-balanced weights 4, no bottleneck 4, the temporal dynamics of ${\bf Q}{\bf Q}^{T}$ are
@@ C.2 Dynamics of the representation from the Lazy to the Rich Regime
The lazy and rich regimes are defined by the dynamics of the NTK of the network. Lazy learning occurs when the NTK is constant, rich learning occurs when it is not. (Farrell et al. (2023))
@@ C.2.1 Lambda-balanced singular value
Under a $\lambda$-Balanced initialization 4, if the network function $\bm{W}_{2}\bm{W}_{1}(t)=\bm{U}(t)\bm{S}(t)\bm{V}^{T}(t)$ is full rank and we define $\bm{S_{\lambda}}(t)=\sqrt{\bm{S}^{2}(t)+\frac{\lambda^{2}}{4}\mathbf{I}}$. , then we can recover the parameters $\bm{W}_{2}(t)=\bm{U}(t)\bm{S}_{2}(t)\bm{R}^{T}(t)$, $\bm{W}_{1}(t)=\bm{R}(t)\bm{S}_{1}(t)\bm{V}^{T}(t)$ up to time-dependent orthogonal transformation $\bm{R}$(t) of size $N_{h}\times N_{h}$, where
@@
We write the initialisation for our setting ${\bf W}_{2}{\bf W}_{1}=\bm{USV}^{T}$. We therefore can write without loss of generality the weight matrices ${\bf W}_{2}=\bm{U}\bm{S}_{2}\bm{G}_{2}$ and ${\bf W}_{1}=\bm{G}_{1}\bm{S}_{1}\bm{V}^{T}$. In this case we requiere that $\bm{G}_{2}=\bm{G}_{1}^{-1}$ and $\bm{S}_{1}\bm{S}_{2}=\bm{S}$.
We assume the balanced property such that $\bm{W}_{2}^{T}\bm{W}_{2}-\bm{W}_{1}\bm{W}_{1}^{T}=\lambda\bm{I}$. We know this holds for any $t$ since this is a conserved quantity in linear networks. The matrices ${\bf W}_{1}{\bf W}_{1}^{T}$ and ${\bf W}_{2}^{T}{\bf W}_{2}$ are symmetric, which consequently implies that their singular vectors are orthogonal. Consequently, in our specific scenario, we require that $\bm{G}_{1}$ and $\bm{G}_{2}$ are orthogonal matrices and it follows that
$\bm{G}_{2}=\bm{G}_{1}^{-1}=\bm{G}_{1}^{T}$ (142)
$\bm{G}_{2}=\bm{G}_{1}^{T}=\bm{R}^{T}$ (143)
We can write $\bm{W}_{2}=\bm{U}\bm{S}_{2}\bm{R}^{T},\bm{W}_{1}=\bm{R}\bm{S}_{1}\bm{V}^{T}$, where $\bm{R}$ is an orthonormal matrix and $\bm{S}_{2},\bm{S}_{1}$ are diagonal (possibly rectangular) matrices.
@@ C.2.4 NTK movement
The theorem above demonstrates that as $|\lambda|\to\infty$, the NTK of a $\lambda$-Balanced network remains constant. This indicates that the network operates in the lazy regime throughout all training steps. The $\lambda$-balanced condition imposes a relationship between the singular values of the two weight matrices. Specifically, if ${\bf W}_{2}$ and ${\bf W}_{1}$ are $\lambda$-balanced and satisfy ${\bf W}_{2}{\bf W}_{1}={\bf\Sigma_{yx}}$, then for arbitrary singular values $a_{i},b_{i},$ and $s_{i}$, the following relations hold:
$a_{i}^{2}-b_{i}^{2}=\lambda,\quad a_{i}\cdot b_{i}=s_{i}.$
As $\lambda$ increases, the value of $b_{i}$ must decrease. In the limit as $\lambda\to\infty$, $a_{i}^{2}\to\lambda$ and $b_{i}^{2}\to 0$. From the first equation, when $b_{i}^{2}\to 0$, $a_{i}^{2}\to\lambda$. Since these equations apply to all singular values of the matrices, it follows that for all $i$, $a_{i}^{2}\to\lambda$, leading to the conclusion that:
${\bf W}_{2}^{T}{\bf W}_{2}=\lambda{\bf I},$
as expected. Consequently, the task representation becomes task-agnostic in ${\bf W}_{1}$. The intuition here is that the weights are constrained by the need to fit the data, which bounds their overall norms. The $\lambda$-balanced condition further specifies a relationship between these norms, and as $|\lambda|$ increases, this constraint tightens, driving ${\bf W}_{2}$ toward the identity matrix. In this regime, the network behaves similarly to a shallow network, with $\lambda$ acting as a toggle between deep and shallow learning dynamics. This finding is significant as it highlights the impact of weight initialization on learning regimes.
@@ C.3 Representation robustness and sensitivity to noise
As derived in (Braun et al., 2025), the expected mean squared error under additive, independent and identically distributed input noise with mean $\mu=0$ and variance $\sigma_{\bf x}^{2}$ is
$\left\langle\frac{1}{2P}\sum_{i=1}^{P}||{\bf W}_{2}{\bf W}_{1}\left({\bf x}_{i}+\xi_{\bf x}\right)-{\bf y}_{i}||_{2}^{2}\right\rangle_{\xi_{\bf x}}=\sigma_{\bf x}^{2}||{\bf W}_{2}{\bf W}_{1}||_{F}^{2}+c,$ (171)
@@
As derived in (Braun et al., 2025), under the assumption of whitened inputs (Assumption 4), in the case of additive parameter noise with $\mu=0$ and variance $\sigma_{\bf W}^{2}$, the expected mean squared error is
@@ C.4 Effect of the architecture from the lazy to the Rich Regime
As $\lambda$ increases, various parts of the network display different learning dynamics: some components adjust rapidly, converging exponentially with $\lambda$, while others adapt more slowly, with their convergence rate inversely proportional to $\lambda$, leading to a slow adaptation.
@@
The first term exhibits exponential decay approaching zero as time progresses:
@@
In the limit as lambda gets large the rate of learning is given by
$\lim_{\lambda\to\infty}\frac{\sqrt{\lambda^{2}\mathbf{I}+4\tilde{\bm{S}}^{2}}}{2}=\frac{\lambda\mathbf{I}}{2},$ (178)
The second term also decays over time
@@
but in the limit as lambda gets large the rate of learning is given by
$\lim_{\lambda\to\infty}\frac{\sqrt{\lambda^{2}\mathbf{I}+4\tilde{\bm{S}}^{2}}-\lambda\mathbf{I}}{2}=\frac{\tilde{\bm{S}}^{2}}{\lambda}$ (179)
Thus, as $\lambda$ increases, the convergence rate slows for certain parts of the network, while other components continue to learn more quickly. This explains the delay observed in the delayed-rich regime. ∎
@@ D.3 Appendix: Generalization and structured learning
D.4 Appendix: Finetuning
It is a common practice to pretrain neural networks on a large auxiliary task before fine-tuning them on a downstream task with limited samples. Despite the widespread use of this approach, the dynamics and outcomes of this method remain poorly understood. In our study, we provide a theoretical foundation for the empirical success of fine-tuning, aiming to improve our understanding of how performance depends on the initialization. We’re interested in understanding how changing the $\lambda$-balancedness after pre-training may impact fine-tuning on a new dataset. We use $\lambda_{PT}$ to denote how networks are first initialized prior to pretraining, and $\lambda_{FT}$ to how they are re-balanced after pre-training and before fine-tuning on a new task. Similar to the previous section, we first train each network (for different values of $-10\leq\lambda_{PT}\leq 10$) on the hierarchical semantic learning task. We then change the $\lambda$-balancedness of each network (for different values of $-10\leq\lambda_{FT}\leq 10$) and retrain on a new dataset to observe how this impacts fine-tuning for different values and compare to networks that are not re-balanced to some $\lambda_{FT}$ ($\lambda_{FT}=\emptyset$) after initial pre-training.
In particular, to reset the $\lambda$-balancedness of a pretrained network to $\lambda_{FT}$, we rescale the singular values of each layer ($\bm{S}_{1},\bm{S}_{2}$) using the singular values of the entire network function ($\bm{S}=\bm{U}^{T}{\bf W}_{2}{\bf W}_{1}\bm{V}$), while keeping the left and right singular vectors of the network unchanged.
We consider three different tasks to fine-tune the networks on. In the first, we add an existing feature from one item to another item in the hierarchy in order to disrupt the structure of the left and right singular vectors. In the second task, we consider the same reversal learning task discussed in Appendix D.2, where one column of the right singular vectors are reversed such that $-\bm{v}_{i}=\tilde{\bm{v}}_{i}$. Finally, we consider a scaled version of the hierarchy where each singular value is scaled by 2.
Across all the tasks we consider, we consistently find that fine-tuning performance improves as networks are re-balanced to larger values of $\lambda_{FT}$ and, conversely, decreases as $\lambda_{FT}$ approaches 0. Networks re-balanced to $\lambda_{FT}=0$ also learn more slowly compared to $|\lambda_{FT}|>0$. Interestingly, when studying networks that are not re-balanced prior to finetuning ($\lambda_{FT}=\emptyset$; but are first initialized prior to pretraining to $\lambda_{PT}$), we see that they perform similarly on the new tasks to networks that are re-balanced to $\lambda_{FT}=\lambda_{PT}$.
Figure 14: Fine-tuning performance on three tasks for different re-balancing $\lambda_{FT}$. A After training on the hierarchical semantic learning task (Section 5), networks are re-balanced and trained on one of three tasks: adding an existing feature from one item to another item in the hierarchy (left), the reversal learning task in Appendix D.2 (center), or a scaled version of the hierarchy where each singular value is scaled by 2 (right). B Change in loss on the new task across different $\lambda_{FT}$ for different $\lambda_{PT}$. As $\lambda_{FT}$ approaches 0, the loss on the new task increases across all $\lambda_{PT}$. Interestingly, networks that are not re-balanced prior to fine-tuning ($\lambda_{FT}=\emptyset$) perform similarly to networks that are re-balanced to the same values ($\lambda_{FT}=\lambda_{PT}$). C Dynamics of the loss across the first pre-training task and the new fine-tuning task. Networks re-balanced to $\lambda_{FT}=0$ consistently learn slower across all tasks compared to networks that are re-balanced to larger magnitude values ($|\lambda_{FT}|>0$
In this work, we derive the precise dynamics of two-layer linear. While straightforward in design, these architectures are foundational in numerous machine learning applications, particularly in the implementation of Low Rank Adapters (LoRA)Hu et al. (2022). A key innovation in LoRA is to parameterize the update of a large weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ within a language model as ΔW = $\Delta\textbf{W = AB}$, the product of two low-rank matrices ${\bf A}\in\mathbb{R}^{d\times r}$ and ${\bf B}\in\mathbb{R}^{r\times d}$, where only ${\bf A}$ and ${\bf B}$ are trained. To ensure $\Delta{\bf W}=0$ at initialization, it is standard practice to initialize ${\bf A}\sim\mathcal{N}(0,\sigma^{2})$ and ${\bf B}=0$ ( Hu et al. (2022); Hayou et al. (2024). It is noteworthy that this parameterization, $\Delta{\bf W=AB}$, effectively embeds a two-layer linear network within the language model. When $r\ll d$, this initialization scheme approximately adheres to our $\lambda$-balanced condition, with $\sigma^{2}$ playing the role of the balance parameter $\lambda$. Investigating how the initialization scale of ${\bf A}$ and ${\bf B}$ influences fine-tuning dynamics under LoRA, and connecting this to our work on $\lambda$-balanced two-layer linear networks and their role in feature learning, represents an intriguing avenue for future exploration. This perspective aligns with recent studies suggesting that low-rank fine-tuning operates in a “lazy” regime, as well as work examining how the initialization of ${\bf A}$ or ${\bf B}$ affects fine-tuning performance Malladi et al. (2023); Hayou et al. (2024). Our framework offers a potential bridge to understanding these phenomena more comprehensively. While a detailed exploration of fine-tuning performance lies beyond the scope of this work, it remains an important direction for future research.